OpenStudy (anonymous):

X1 and X2 are independently distributed random variables with P(X1=Ɵ+1) = P(X1=Ɵ-1) = 1/2 P(X2=Ɵ-2) = P(X2=Ɵ+2) = 1/2 i)Find the values of a and b which minimize the variance of Y=aX1 + bX2 subject to the condition that E[Y]=Ɵ. ii)What is the minimum value of this variance? The answers at the back of the book says that a and b is 4/5 and 1/5 respectively and the variance is 4/5. I don't know how they got that answer and I'm not even sure where to start...

OpenStudy (anonymous):

check this I worked out that E[X1]=E[X2]=θ and then i got var(X1)=1 and var(X2) = 4 using var(X) = E[X^2] - (E[X])^2 by equating E[Y]=θ=E[aX1+bX2] i get a+b=1

OpenStudy (anonymous):

don't know what to do next

OpenStudy (anonymous):

Yes, i to the exact same and im stuck on that same point to...

OpenStudy (phi):

Var(aX+bY)= a^2 Var(X) + b^2 Var(Y) + 2ab Cov(XY) independently distributed random variables means Cov(XY)=0 so for this problem Var(Y)= 1*a^2 + 4*b^2 also, as noted above, a+b=1 so a= 1-b. substitute into the equation for Var(Y): Var(Y)= (1-b)^2 + 4b^2 simplify, take the derivative wrt b, set to zero and solve for b to find the value of b that minimizes Var(Y)

OpenStudy (anonymous):

thanks so much for the help!^-^